# Combinatorial Selection with Costly Information

### v1

@@ 7.2 Failure of Local Approximation for PBOI with Rewards
For all $\alpha>1/2$, there exists a box that does not admit a local $\alpha$-approximation.
Proof.
For any $n\gg 1$, consider the box $\mathcal{B}^{n}$ where
$\mathcal{D}=\begin{cases}1&\text{w.p. }1-\frac{1}{n^{2}}\\ n^{3}&\text{w.p. }\frac{1}{n^{2}}\\ \end{cases}$
and $c=n-1$. Note that the mean of this distribution, $m$, is equal to $n+1-\frac{1}{n^{2}}$, and the Gittins index $g$ is $n^{2}$. We get that $\mathbf{E}[\max\{\alpha W^{*},0\}]=\alpha m$ and $\mathbf{E}[\max\{W^{\mathrm{open}},0\}]=m-c$. Similarly, if $y=m$ then
$\mathbf{E}[\max\{\alpha W^{*},y\}]=\max\{\mathbf{E}[\max\{\alpha W^{\mathrm{open}},m\}],\alpha m\}=\mathbf{E}[\max\{\alpha W^{\mathrm{open}},m\}]$
and then a straightforward calculation gets us that for all $\alpha\geq\frac{1}{2}$ and sufficiently large $n$,
$\mathbf{E}[\max\{\alpha W^{\mathrm{open}},m\}]=\mathbf{E}_{X\sim\mathcal{D}}[\max\{\alpha X,m\}]-\alpha c=\alpha+m\left(1-\frac{1}{n^{2}}\right).$
Thus, if we want to satisfy
$(1-p)\mathbf{E}[\max\{W^{\mathrm{open}},y\}]+p\max\{m,y\}\geq\mathbf{E}[\max\{\alpha W^{*},y\}]$
at $y=0$, we can plug in the quantities computed above and reorganize to get the bound $p\geq\frac{m(\alpha-1)}{c}+1$. Similarly, we can plug in our computed quantities when $y=m$ to get that when $n$ is sufficiently large and $\alpha\geq\frac{1}{2}$,
$p\leq\frac{\alpha-1}{\frac{m}{n^{2}}-1}.$
Putting both of these bounds together, we get the following bound on $\alpha$,
$\alpha\leq\frac{m(c-m+n^{2})}{mn^{2}+cn^{2}-m^{2}},$
which goes to $\frac{1}{2}$ from above as $n\to\infty$. Therefore, for any $\alpha>\frac{1}{2}$, we can choose $n$ large enough that $\mathcal{B}^{n}$ does not admit a local $\alpha$-approximation. ∎
To understand the behavior that leads to the worst case local $\alpha$-approximations, it will be illustrative to consider a box $\mathcal{B}^{n}$ for large $n$ as defined in the proof of Theorem 7.2. Figure 7 plots the value of grabbing versus inspecting as a function of the outside value $y$ for the box $\mathcal{B}^{20}$. As we can see, when $y$ is small we have a strong preference for grabbing the box, and when $y$ is larger we have a very slight preference to inspect the box. Looking at this plot makes it intuitively clear that the “always grab” policy performs quite well, since in the worst case there is only a small additive sub-optimality factor. However, local $\alpha$-approximation requires a small multiplicative sub-optimality factor, which, as the above proof demonstrates, the “always grab” policy does not achieve.
Figure 7: The red curve, $\max\{m,y\}$, represents the value from grabbing the box $\mathcal{B}^{20}$ (as defined in the proof of Theorem 7.2). The blue curve, $\mathbf{E}[\max\{W^{\mathrm{open}},y\}]$, represents the value from inspecting the box $\mathcal{B}^{20}$. This box admits at most a local $0.537$-approximation. By considering $\mathcal{B}^{n}$ as $n\to\infty$, Theorem 7.2 shows that for all $\varepsilon>0$, there exist boxes admit at most a local $(1/2+\varepsilon)$-approximation.
To capture this intuition, in Section 7.3, we introduce semilocal $(\alpha,\beta)$-approximation (Definition 11), where $\alpha$ captures the multiplicative suboptimality while $\beta$ captures the additive suboptimality. We show that semilocal $(\alpha,\beta)$-approximations compose in the matroid selection setting leading to an overall $(\alpha-\beta)$-approximation (Theorem 7.3). Then, in Section 7.4, we show that all boxes admit a semilocal $(0.682,0.1)$-approximation (Theorem 7.7), guaranteeing a $0.582$-approximation in the matroid selection setting and thus breaking the existing barrier of $1/2$.
7.3 Semilocal Approximations and Their Composition
Definition 11 (Semilocal Approximation).
Consider a box $\mathcal{B}=(\mathcal{D},c)$. We say that a probability $p\in[0,1]$ is a semilocal $(\alpha,\beta)$-approximation for $\mathcal{B}$ if for all $y$,
$(1-p)\mathbf{E}[\max\{W^{\mathrm{open}},y\}]+p\max\{m,y\}\geq\mathbf{E}[\max\{\alpha W^{*},y\}]-p\beta m.$ (3)
If there exists a hedging probability that is a semilocal $(\alpha,\beta)$-approximation for the box $\mathcal{B}$, then we say $\mathcal{B}$ admits semilocal $(\alpha,\beta)$-approximation.
Rewriting 3 as
$(1-p)\mathbf{E}[\max\{W^{\mathrm{open}},y\}]+p(\max\{m,y\}+\beta m)\geq\mathbf{E}[\max\{\alpha W^{*},y\}].$
makes it clear that semilocal $(\alpha,\beta)$-approximation is basically a local $\alpha$-approximation where we have boosted the value attained from grabbing by an additive factor proportional to the mean. It is important that we do this only for grabbing—boosting the value of inspecting by an additive factor would lead to trivial guarantees when we compose semilocal approximations in the matroid selection setting.
Input : A matroid PBOI instance $\mathbb{I}=(\mathcal{B}_{1},\dots,\mathcal{B}_{n},\mathcal{F})$ (normalized as described in 1)
A vector of probabilities $(p_{1},\dots,p_{n})$
1 Relabel the boxes such that $m_{1}\geq\dots\geq m_{n}$
@@ 3       If $K_{i}=1\textbf{ and }L_{i}=1$ :
8 Mark the boxes in $S^{\mathrm{grab}}$ as “grab” and the rest as “open”
9 Run the frugal Gittins policy (Singla, 2017) on the resulting mandatory-inspection instance
@@
Theorem 7.3 (Composition of Semilocal Approximations).
Let $\alpha>\beta\geq 0$, let $\mathbb{I}=(\mathcal{B}_{1},\dots,\mathcal{B}_{n},\mathcal{F})$ be a max-matroid-CICS instance, and let $(p_{1},\dots,p_{n})$ be a vector of probabilities such that $p_{i}$ is a semilocal $(\alpha,\beta)$-approximation for $\mathcal{B}_{i}$. Then there exists a committing policy, namely Algorithm 3, that is an $(\alpha-\beta)$-approximation for the CICS instance $\mathbb{I}$.
@@
There are two main steps of the proof, each stated and proved in a lemma below. We express both steps in terms of the random variables
$W^{\mathrm{alg}}_{i}=\begin{cases}W^{\mathrm{open}}_{i}&\text{if }K_{i}L_{i}=0\\ m_{i}&\text{if }K_{i}L_{i}=1,\end{cases}$
where $K_{i}$ and $L_{i}$ are as defined in Algorithm 3. One can think of $W^{\mathrm{alg}}_{i}$ as the surrogate value of box $i$ conditional on the state of the algorithm at Algorithm 3.
  1. The first step, Lemma 7.4, is to express the value achieved by Algorithm 3 in terms of $W^{\mathrm{alg}}_{i}$.
  2. The second step, Lemma 7.5, is to compare the resulting expression to an upper bound on the optimal value. This step uses the semilocal $(\alpha,\beta)$-approximation guarantee from Definition 11, which gives us a relationship between $W^{\mathrm{alg}}_{i}$ and $W^{*}_{i}$.
Combining the lemmas yields
[…]
  1. Theorem E.4 implies
    $\mathbf{E}\bigg{[}\max_{S\in\mathcal{F}}\sum_{i\in S}\alpha W^{*}_{i}\bigg{]}\geq\alpha\mathbf{E}[\text{value of optimal policy}].$
  2. Because Algorithm 3 ensures $S^{\mathrm{grab}}\in\mathcal{F}$ by construction, the following policy is feasible: “Compute $S^{\mathrm{grab}}$ as in Algorithm 3, but then simply grab the boxes in $S^{\mathrm{grab}}$.” This algorithm achieves value $\mathbf{E}\big{[}\sum_{i\in S^{\mathrm{grab}}}m_{i}\big{]}$, which means
[…]
Lemma 7.4.
[…]

### v2

@@ 8.3 Semilocal Approximation
Let $\alpha>\beta\geq 0$, and let $\mathbb{I}=(\mathcal{B}_{1},\dots,\mathcal{B}_{n},\mathcal{F})$ be any instance of matroid-max-PBOI where each constituent box $\mathcal{B}_{i}$ admits an $(\alpha,\beta)$-semilocal approximation under some probability $p_{i}\in[0,1]$. Then, $\mathrm{CG}(\mathbb{I})\geq\alpha-\beta$.
Input : A normalized matroid-max-PBOI instance $\mathbb{I}=(\mathcal{B}_{1},\dots,\mathcal{B}_{n},\mathcal{F})$.
A vector of probabilities $(p_{1},\dots,p_{n})$.
1 Relabel the boxes such that $\mu_{1}\geq\dots\geq\mu_{n}$
@@ 4       If $K_{i}=1\textbf{ and }L_{i}=1$ :
8
9
10Commit to grabbing the boxes $S^{\mathrm{grab}}$ and opening the boxes in $[n]\setminus S^{\mathrm{grab}}$.
11 Run the optimal (index) policy under the resulting commitment.
@@
The committing policy achieving the bound of Theorem 8.2 is presented in Algorithm 2. Importantly, this policy does not commit to grabbing each box $\mathcal{B}_{i}$ with probability $p_{i}$ independently (as would be the case when composing local approximation guarantees). Instead, it examines the boxes in decreasing mean order, flips a coin with probability $p_{i}$ for each box $\mathcal{B}_{i}$ and only commits to grabbing it if (i) the coin flip succeeds and (ii) the set of boxes that are committed to the grab action after inserting $\mathcal{B}_{i}$ remains an independent set of the underlying matroid. Once the commitments for each box have been specified, we simply run the optimal policy for the resulting matroid-max-CICS instance over Markov chains (equivalently, we substitute each grab box $(\mathcal{D},c)$ with a box $(\mu,0)$ and run Weitzman’s algorithm for the mandatory inspection setting). Therefore, as long as the probabilities $p_{i}$ achieving the semilocal approximation can be efficiently computed, the entire algorithm runs in polynomial time. The proof of Theorem 8.2 is deferred to Appendix I.
We note that both the definition and composition algorithm for semilocal approximation are specifically tailored to the PBOI model, though we believe modest extensions beyond PBOI may be possible. The main limitaiton is that the proof of Theorem 8.2 relies crucially on the fact that if we choose to grab a box, then the box’s terminal value is deterministic, namely $\mu$. It is thus natural to try extending semilocal approximation to other Pandora’s box variants that have a grab-like action that yields a deterministic terminal value, but we leave this to future work.
8.4 Breaking the 0.5 Barrier
We are finally ready to prove Theorem 8.1. From Theorem 8.2, it suffices to prove that all optional inspection boxes admit an $(\alpha,\beta)$-semilocal approximation for some $\alpha,\beta$ such that $\alpha-\beta\geq 0.582$. This is achieved by the following lemma, proven in Appendix I.
Lemma 8.3.
Let $\mathcal{B}$ be any (non-degenerate) box with opening cost $c$ and mean value $\mu$. For any $\beta\geq 0$, let
$\alpha(\beta):=\begin{cases}\frac{1}{1+\frac{c}{\mu}-\beta\frac{c}{\mu-c}}&\;\;,\text{if }\frac{1}{1+\frac{c}{\mu}-\beta\frac{c}{\mu-c}}\in(0,1].\\ 1&\;\;,\textnormal{otherwise.}\end{cases}$
Then, the probability $p=\frac{c}{\mu}\alpha(\beta)$ achieves an $(\alpha(\beta),\beta)$-semilocal approximation for $\mathcal{B}$.
By optimizing over $\alpha(\beta)-\beta$, we instantiate Lemma 8.3 with $\beta=0.1$, to obtain the corresponding probability. It is straightforward to verify (using a computer algebra system or numerical solver) that since $\frac{c}{\mu}\in[0,1]$ (recall that we have normalized all degenerate boxes), we have
$\left(1+\frac{c}{\mu}-\frac{c}{10(\mu-c)}\right)^{-1}\geq 0.682$
whenever $\big{(}1+\frac{c}{\mu}-\frac{c}{10(\mu-c)}\big{)}^{-1}\in(0,1]$. This means $\alpha(0.1)\geq 0.682$, and thus all boxes admit a $(0.682,0.1)$-semilocal approximation, from which the proof of Theorem 8.1 follows.
Critical to the above argument is that Lemma 8.3 implies there exists a universal pair $(\alpha,\beta)$, namely $(0.682,0.1)$, such that all boxes admit an $(\alpha,\beta)$-semilocal approximation. It would not suffice to show only that for each box, there exists a box-specific pair $(\alpha^{\prime},\beta^{\prime})$ such that the box admits a $(\alpha^{\prime},\beta^{\prime})$-semilocal approximation, even if we always had $\alpha^{\prime}-\beta^{\prime}\geq 0.582$. This is because Theorem 8.2 requires all boxes to admit a $(\alpha,\beta)$-approximation for the same pair $(\alpha,\beta)$.
Appendix
Appendix A The Maximization Setting
A.1 Amortization for Markov Chains
The cost amortization of a Markov chain is defined in the same manner, with the difference that instead of increasing the terminal values of the trajectories to obtain surrogate costs, we now decrease them to obtain surrogate values. Furthermore, the index of a state now corresponds to the maximum surrogate value among all downwards trajectories instead of the minimum. Formally:
Definition 13 (Cost Amortization for Maximization).
A cost amortization for any Markov chain $\mathcal{M}=(S,\sigma,A,c,\mathcal{D},V,T)$ is a non-negative vector $b=\{b_{s\tau}\}_{s\in S,\tau\in\Upsilon(s)}$ satisfying $\sum_{\tau\in\Upsilon(s)}p(\tau)b_{s\tau}=p(s)c(s)$ for all states $s\in S$. Based on this amortization, we define:
  • The amortized value of a trajectory $\tau\in\Upsilon$ as $\rho_{b}(\tau):=v(\tau)-\sum_{s\in\tau}b_{s\tau_{s}}$.
  • The surrogate value of the Markov chain $\mathcal{M}$ as the random variable $\rho_{\mathcal{M},b}$ that takes on value $\rho_{b}(\tau)$ for $\tau\in\Upsilon$ with probability $p(\tau)$.
  • The index of a state $s\in S$ of the Markov chain $\mathcal{M}$ as $I_{\mathcal{M},b}(s):=\max_{\tau\in\Upsilon:s\in\tau}\rho_{b}(\tau)$.
Lemma A.1 (Markov chain upper upper for maximization settings).
Consider any instance $\mathbb{I}=(\mathbb{M},\mathcal{F})$ of matroid-max-CICS over Markov chains and let $b_{i}$ be any cost amortization of $\mathcal{M}_{i}$ with surrogate value $\rho_{i}:=\rho_{\mathcal{M}_{i},b_{i}}$ for all $i\in[n]$. Then, $\mathrm{OPT}(\mathbb{I})\leq\operatorname{E}\left[\max_{S\in\mathcal{F}}\sum_{i\in S}\rho_{i}\right]$.
Up next, we extend our definition of water filling amortization to the maximization setting. Like before, the water draining cost amortization is described algorithmically in a bottom-up fashion where states are considered in the chain in reverse topological order, starting from the terminals up towards the root $\sigma$. Trajectories $\tau\in\Upsilon(t)$ for terminal states $t\in T$ are singletons and do not carry cost shares. Consider a state $s$ such that the cost shares for all states reachable from it have been determined. The state distributes its total cost $c(s)$ across its downstream trajectories $\tau\in\Upsilon(s)$, starting from those with the maximum current value, until the equation $\sum_{\tau\in\Upsilon(s)}p(\tau)b_{s\tau}=p(s)c(s)$ is satisfied. In other words, instead of increasing the cost of the minimum-cost trajectory, we now decrease the value of the maximum-value trajectory. We use $W^{*}_{\mathcal{M}}$ to denote the water draining surrogate value of a Markov chain $\mathcal{M}$, and $I^{*}_{\mathcal{M}}(s)$ to denote the water draining index of a state $s$ in $\mathcal{M}$.
Definition 14 (Water Draining Index policy).
The water draining index policy for an instance $\mathbb{I}=(\mathbb{M},\mathcal{F})$ of matroid-max-CICS over Markov chains selects at every step the Markov chain $i^{*}=\operatorname{argmax}_{i\in\mathcal{F}_{S}}I^{*}_{i}(s_{i})$, where $s_{i}$ is the current state of Markov chain $\mathcal{M}_{i}$; $S$ is the set of terminated (selected) Markov chains; and $\mathcal{F}_{S}=\{i:S\cup\{i\}\in\mathcal{F}\}$.
From the same observations as in the proof of Lemma 3.1, it immediately follows that this algorithm achieves two desired properties: it always selects the feasible set of Markov chains with maximum total surrogate value, and whenever it advances a state $s\in S$ that contributes to the surrogate value of a downwards trajectory $\tau\in\Upsilon(s)$ (i.e. $b_{s\tau}>0$) and nature realizes $\tau$, the algorithm will select it. From this, and the fact that the algorithm that greedily adds the maximum weight feasible element is optimal for the matroid independent set problem, the following counterpart to Theorem 3.2 follows, establishing the optimality of the Water Draining Index policy for matroid-max-CICS.
Theorem A.2.
For any matroid-max-CICS instance $\mathbb{I}=(\mathbb{M},\mathcal{F})$ over Markov chains, the expected utility of the water draining index policy is equal to $\operatorname{E}\left[\max_{S\in\mathcal{F}}\sum_{i\in S}W^{*}_{\mathcal{M}_{i}}\right]$. The policy is therefore optimal for $\mathbb{I}$.
A.2 Optimality Curves
The definition of a local game $(\mathcal{M},y)$ naturally extends to the maximization setting; at any step, the decision maker can either advance the Markov chain $\mathcal{M}$ at a cost (until it reaches a terminal state in which case it may collect its value and terminate), or collect the reward $y$ of the outside option and terminate. Like before, we use $f_{\mathcal{M}}(y)$ to denote the utility of the optimal adaptive policy in this game. From Theorem A.2, we immediately have that
Definition 15 (Water draining surrogate values for MDPs).
Let $\mathcal{M}$ be an MDP with optimality curve $f_{\mathcal{M}}$. The water draining surrogate value for $\mathcal{M}$ is the random variable $W^{*}_{\mathcal{M}}$ generated from the CDF $\frac{d}{dy}f_{\mathcal{M}}(y)$. That is, $W^{*}_{\mathcal{M}}$ is the random variable satisfying $f_{\mathcal{M}}(y)=\operatorname{E}\left[\max(y,W^{*}_{\mathcal{M}})\right]$ for all $y\in\real$.
A.3 Amortization for MDPs
Lemma A.3.
For any MDP $\mathcal{M}=(S,\sigma,A,c,\mathcal{D},V,T)$ and any deterministic commitment $\pi\in\mathcal{C}\left(\mathcal{M}\right)$, generating a Markov chain $\mathcal{M}^{\pi}$ with states $S_{\pi}\subseteq S$, realizable trajectories $\Upsilon_{\pi}\subseteq\Upsilon$ and a distribution $p_{\pi}$ over trajectories and reachable states, there exists an amortized cost function $\rho^{\pi}:\Upsilon_{\pi}\mapsto\Delta(\real)$ mapping trajectories to distributions over costs and a non-negative cost sharing vector $b^{\pi}=\{b^{\pi}_{s\tau}\}_{s\in S_{\pi},\tau\in\Upsilon_{\pi}(s)}$, such that the following properties hold:
  1. 1.
    Cost Sharing: the amortized cost of a trajectory pays for its own acceptance value and the cost shares it sends to upstream states. For all $\tau\in\Upsilon_{\pi}$, it holds that $\operatorname{E}\left[\rho^{\pi}(\tau)\right]=v(\tau)-\sum_{s\in\tau}b^{\pi}_{s\tau_{s}}$.
  2. 2.
    Cost Dominance: the cost shares received by any state pay towards its action cost, but do not overpay. For all $s\in S_{\pi}$, it holds that $\sum_{\tau\in\Upsilon_{\pi}(s)}p_{\pi}(\tau)b^{\pi}_{s\tau}\leq p_{\pi}(s)c(\pi(s))$.
  3. 3.
    Action Independence: sampling a trajectory $\tau\sim p_{\pi}$ and then sampling from the amortized cost distribution $\rho^{\pi}(\tau)$ generates a random surrogate cost for the MDP. This random variable is distributed identically to the water draining surrogate cost $W^{*}_{\mathcal{M}}$.
@@ A.3 Amortization for MDPs
and by defining $\hat{Z}_{j}:=\min\{g_{j},Z_{j}\}$, we have
which in turn allows us to argue that
for all $y\in\real$ and $j\in[k]$.
  1. 1.
  2. 2.
We can now define the amortized cost functions $\rho^{\pi}$ and the cost sharing vectors $b^{\pi}$. Fix any deterministic commitment $\pi\in\mathcal{C}\left(\mathcal{M}\right)$ and let $j=\pi(\sigma)\in[k]$ be the fixed action that $\pi$ takes at state $\sigma$. For each state $s\in R_{j}$, we use $\pi|s\in\mathcal{C}\left(\mathcal{M}_{s}\right)$ to denote the commitment $\pi$ after we transition to state $s$; note that this is also a deterministic commitment. By definition, each MDP $\mathcal{M}_{s}$ has horizon up to $H$ and thus by the induction hypothesis, it admits a mapping $\rho^{\pi|\sigma}(\cdot)$ and a non-negative cost sharing vector $b^{\pi|\sigma}$, satisfying the properties of the lemma. Now, consider any trajectory $\tau\in\Upsilon_{\pi}$ and observe that $\tau=\{\sigma,\tau_{s}\}$ for some $s\in R_{j}$ and some $\tau_{s}\in\Upsilon_{\pi}(s)$. We define
$\rho^{\pi}(\tau):=m_{j}(\min\{g_{j},\rho^{\pi|s}(\tau_{s})\})$
and
$b^{\pi}_{\sigma\tau}:=\operatorname{E}\left[\rho(\tau)\right]-\operatorname{E}\left[\rho^{\pi|s}(\tau_{s})\right]$
and for all other $s\in S_{\pi}\setminus\{\sigma\}$ and $\tau\in\Upsilon_{\pi}(s)$, we use the same cost share $b^{\pi}_{s\tau}=b^{\pi|s}_{s\tau}$ that was used in $\mathcal{M}_{s}$. The proof of the three properties follows precisely the same steps as Lemma 3.5. ∎
From Lemma A.3, we immediately obtain the following counterpart of Theorem 3.4; the proof follows exactly the same steps as the proof of Theorem 3.4 that we presented in Appendix D, and is thus omitted.
Theorem A.4.
In any instance $\mathbb{I}=(\mathbb{M},\mathcal{F})$ of matroid-max-CICS, $\mathrm{OPT}(\mathbb{I})\leq\operatorname{E}\left[\max_{S\in\mathcal{F}}\sum_{i\in S}W^{*}_{\mathcal{M}_{i}}\right]$.
A.4 Local Approximation
Definition 16 (Local approximation for maximization settings).
Let $\mathcal{M}$ be any MDP and let $\alpha\in(0,1]$. We say that a commitment $\pi\in\mathcal{C}\left(\mathcal{M}\right)$ is an $\alpha$-local approximation for $\mathcal{M}$ if $f_{\mathcal{M}^{\pi}}(\alpha y)\geq\alpha\cdot f_{\mathcal{M}}(y)$ for all $y\in\real$.
Theorem A.5 (Composition theorem for maximization settings).
Let $\mathbb{I}=(\mathbb{M},\mathcal{F})$ be any instance of matroid-max-CICS, where each constituent MDP $\mathcal{M}_{i}$ admits an $\alpha$-local approximation under some commitment $\pi_{i}\in\mathcal{C}\left(\mathcal{M}_{i}\right)$. Then, $\mathrm{CG}(\mathbb{I})\geq\alpha$.
By combining Theorem A.5 with the optimality of the Water Draining Index policy (Theorem A.2) and the upper bound of Theorem A.4, we obtain a way to efficiently approximate the optimal solution for matroid-max-CICS assuming that the underlying MDPs achieve good local approximation guarantees.
Appendix B The Combinatorial Setting
In this section, we describe how our entire framework extends to combinatorial settings beyond the case of matroids. We will state our results in full generality and distinguish between minimization and maximization whenever needed. The framework we consider is based on (Singla, 2017) and its followup (Gupta et al., 2019), where the authors develop a technique for combinatorial selection over Markov chains. In this section, we show that under local approximation, their results can be seamlessly extended to arbitrary MDPs.
Definition 17 (CICS).
$\sum_{t\in T}v(t)+h(S)+\sum_{a\in A}c(a)$
$\sum_{t\in T}v(t)+h(S)-\sum_{a\in A}c(a).$
Theorem B.1.
Let $\mathbb{I}=(\mathcal{M}_{1},\cdots,\mathcal{M}_{n},\mathcal{F},h)$ be any instance of CICS. For each $i\in[n]$, let $W^{*}_{\mathcal{M}_{i}}$ be the water filling (resp. water draining) surrogate costs in the minimization (resp. maximization) setting. Then:
  1. 1.
    For the minimization setting, $\mathrm{OPT}(\mathbb{I})\geq\operatorname{E}\left[\min_{S\in\mathcal{F}}(\sum_{i\in S}W^{*}_{\mathcal{M}_{i}}+h(S))\right]$.
  2. 2.
    For the maximization setting, $\mathrm{OPT}(\mathbb{I})\leq\operatorname{E}\left[\max_{S\in\mathcal{F}}(\sum_{i\in S}W^{*}_{\mathcal{M}_{i}}+h(S))\right]$.
Theorem B.2.
Let $\mathbb{I}=(\mathcal{M}_{1},\cdots,\mathcal{M}_{n},\mathcal{F},h)$ be any instance of CICS. For each $i\in[n]$, let $W^{*}_{\mathcal{M}_{i}}$ be the water filling (resp. water draining) surrogate costs in the minimization (resp. maximization) setting. Finally, for each $i\in[n]$, let $\pi_{i}\in\mathcal{C}\left(\mathcal{M}_{i}\right)$ be some commitment that $\alpha$-locally approximates $\mathcal{M}_{i}$. Then:
  1. 1.
  2. 2.
By combining Theorem B.1 with Theorem B.2, and assuming that our MDPs admit local approximation guarantees, we are left with the task of optimizing over the CICS instance that is generated by the commitments; observe that this is now an instance over Markov chains. In the case of matroid constraints, we showed that we can always efficiently achieve this via the water filling/draining index policy. However, depending on the combinatorial constraint, efficient optimization might not be possible – consider for example the case where each MDP is a single terminal state corresponding to a set of elements and we need to select a minimum cost set cover.
The final key to the puzzle will be a way of efficiently approximating the optimal policy for a CICS instance $\mathbb{I}=(\mathcal{M}_{1},\cdots,\mathcal{M}_{n},\mathcal{F},h)$ over Markov chains. This is precisely the setting that is considered by Gupta et al. (2019). In this work, it is shown that a sufficient condition to get an efficient $\beta$-approximation to the optimal policy for a CICS instance over Markov chains is for the underlying pair $(\mathcal{F},h)$ to admit a $\beta$-approximate frugal algorithm. We defer the reader to Gupta et al. (2019) for the full details. Here, we will just mention some examples of such combinatorial settings:
  • Matching constraints in the maximization setting admit a $0.5$-approximate frugal algorithm.
Corollary B.3.
  1. 1.
    Each MDP $\mathcal{M}_{i}$ admits an $\alpha$-local approximation.
  2. 2.
Then, $\mathrm{CG}(\mathbb{I})\leq\alpha\cdot\beta$ for minimization, and $\mathrm{CG}(\mathbb{I})\geq\alpha\cdot\beta$ for maximization.
In other words, if (i) we can efficiently compute an $\alpha$-local approximate commitment for each MDP and (ii) the combinatorial setting admits a frugal algorithm, then we can efficiently construct a committing policy that $(\alpha\beta)$-approximates the optimal (non-committing) policy. We are left with the task of proving our extended Theorems B.1 and B.2. It is not hard to see that both proofs are identical to their matroid counterparts and are thus omitted. In particular:
  1. 1.
    The proof of Theorem B.1 follows exactly the same steps as the proofs of Theorem 3.4 (for the minimization setting) and Theorem A.4 (for the maximization setting) that were presented in Appendix D and Appendix A respectively. In these proofs, we separately bounded the cost/utility contribution of each MDP by the corresponding surrogate costs/values and simply invoked the feasibility of the optimal adaptive policy at the end. Thus, the exact same proofs imply Theorem B.1 for any feasibility constraint $\mathcal{F}$ and any cost/reward function $h(\cdot)$.
  2. 2.
    The proof of Theorem B.2 follows exactly the same steps as the proofs of Theorem 4.1 (for the minimization setting) and Theorem A.5 that were presented in Appendix E and Appendix A respectively. In particular, none of these proofs used the fact that $\mathcal{F}$ is a matroid constraint at any point, and it is also straightforward to see that they immediately extend for any cost/reward function $h(\cdot)$.
Appendix C Challenges of Composing Local Conditions
In this section, we demonstrate the challenges of extending Whittle’s condition to obtain bounds on the commitment gap. Fix any instance $\mathbb{I}=(\mathbb{M},\mathcal{F})$ of CICS. Whittle’s condition states that if the (local) commitments $\pi_{i}$ are optimal for all local games $(\mathcal{M}_{i},y)$, then $\Pi=(\pi_{1},\cdots,\pi_{n})$ is optimal for $\mathbb{I}$ and thus the commitment gap is $1$. Naturally, this is a very strong local condition and it is desirable to relax it into an approximation guarantee, while still being able to compose it into a global bound.
A natural approach would be to try and extend this result to approximation guarantees over the local games. However, this does not work. Consider, for example, a minimization MDP $\mathcal{M}$ with two 00-cost actions: the first action leads to a deterministic value of $1$, and the second uniformly leads to a stochastic value of 00 or $1$. Committing to each of these actions leads to an expected cost of $\min(y,1)$ and $1/2\cdot\min(y,1)$ in the local game respectively; the deterministic action achieves a $2$-approximation for all local problems $(\mathcal{M},y)$. Now consider a single-selection CICS with $n$ copies of the above MDP. If we commit to the deterministic action in each MDP, our cost is deterministically $1$, whereas taking the stochastic action in each costs us $1/2^{n}$ – an exponential gap!
A different approach would be to consider the negation of Whittle’s condition: if a (local) commitment $\pi_{i}$ is strictly suboptimal for all local game $(\mathcal{M}_{i},y)$, then no optimal policy for the CICS instance $\mathbb{I}$ will commit to $\pi_{i}$. This could potentially allow us to rule out some of the commitment policies and simplify our instance. However, we show that this isn’t true: there are instances of CICS where the commitment gap is $1$ (and thus the optimal policy is a committing policy) and yet the actions it commits to are unambiguously suboptimal!
The Example.
We consider an instance of single-selection min-CICS consisting of two MDPs ($\mathcal{M}_{1},\mathcal{M}_{2})$. The MDP $\mathcal{M}_{1}$ is a depth $1$ Markov chain, consisting of a single action of cost $3$ and resulting to a uniformly random terminal state of value 00 or $50$. The MDP $\mathcal{M}_{2}$ is of height $2$. On the first level, there are two actions $a_{1},a_{2}$ of cost $c(a_{1})=c(a_{2})=0$. Taking action $a_{1}$ deterministically results to a terminal state of value $5$. Taking action $a_{2}$ results to a terminal state of value 00 with probability $1/2$ and to a non-terminal state $s$ with probability $1/2$. On state $s$, there are two more actions $a_{3},a_{4}$ of cost $c(a_{3})=c(a_{4})=0$. Taking $a_{3}$ results to a terminal of value $12$ and taking $a_{4}$ results to a uniformly random terminal of value 00 or $50$.
Notice that there are three committing policies in $\mathcal{C}\left(\mathcal{M}_{2}\right)$, corresponding to action sequences $\{a_{1}\}$, $\{a_{2},a_{3}\}$ and $\{a_{2},a_{4}\}$; we use $\pi_{1}$, $\pi_{23}$ and $\pi_{24}$ to denote these policies. As we have seen, each of these committing policies $\pi$ implies a Markov chain $\mathcal{M}_{2}^{\pi}$ and an optimality curve for the local game $(\mathcal{M}_{2}^{\pi},y)$, denoting the cost of the optimal policy for the local game as a function of the outside option $y\geq 0$. For this specific example, it is not hard to see that:
  • $f_{1}(y):=f_{\mathcal{M}^{\pi_{1}}_{2}}(y)=\min\{y,5\}$.
  • $f_{23}(y):=f_{\mathcal{M}^{\pi_{23}}_{2}}(y)=0.5\cdot\min\{y,12\}$.
  • $f_{24}(y):=f_{\mathcal{M}^{\pi_{24}}_{2}}(y)=0.25\cdot\min\{y,50\}$.
A pictorial representation of the two MDPs and the corresponding optimality curves for the three committing policies on $\mathcal{M}_{2}$ is shown in Figure 10.
Figure 10: The MDPs $\mathcal{M}_{1}$ (left) and $\mathcal{M}_{2}$ (right). All the actions in $\mathcal{M}_{2}$ have a cost of 00. There are three committing policies for $\mathcal{M}_{2}$ with optimality curves $f_{1}(y)$ (blue), $f_{23}(y)$ (orange) and $f_{24}(y)$ (red).
The optimality curves demonstrate that for any value of $y$, the committing policy $\pi_{23}$ is sub-optimal for the local game $(\mathcal{M}_{2},y)$; in other words, the optimal policy for the local game will never take action $a_{3}$. One could expect that since $a_{3}$ is always suboptimal, the optimal policy would never take this action in the context of the bandit superprocess $(\mathcal{M}_{1},\mathcal{M}_{2})$. However, we can show that this isn’t true. On the contrary, the optimal algorithm for $(\mathcal{M}_{1},\mathcal{M}_{2})$ will commit to $\pi_{23}$!
Claim 3.
The optimal policy for the single-selection min-CICS instance $(\mathcal{M}_{1},\mathcal{M}_{2})$ is the committing policy that commits to the unique commitment for Markov chain $\mathcal{M}_{1}$ and to $\pi_{23}$ for MDP $\mathcal{M}_{2}$.